Physics help please
imagine that youve been invited to try out a new hoversuit, and heres how it works: Someone has set up a large flat sheet, many kilometers across, somewhere on the Earth, and theyve charged the sheet up to a uniform charge density of + P. 3 2x 10-6C/㎡,you are issued a special suit that you wear, and it has controls on it which allow you to charge the sau up to any number of Coulombs (C), positive or negative, that you might want. The you can control the amount of electrical repulsion (or attraction) between the suit charged sheet below idea is that and the you. Assume that the acceleration due to gravity is 10 m/s2, and that your mass is Q. kg. k (a) What is the strength of the electric field in the region above the charged sheet? Answer: XION/C (b) How much charge (and is it positive or negative) must you give the suit if you want to be able to just hover stationary above the charged sheet? Answer:_10 (c) How much charge (and is it positive or negative) must you give the suit if you want to fall up, that is, accelerate upward at the same rate that objects normally accelerate downward due to Earths gravity? Answer:x10C
Physics help please
Answer #1

Electric field of Sheet of Charge ,

E = \sigma/2\epsilon = 3.3x10-6/2*8.854x10-12 = 1.86x105 N/C

To hover , mg = qE => q = mg/E = 69*10/1.86x105 = 3.7 mC (+ve)